# The Cramér-Rao approach and global quantum estimation of bosonic states

### v4

@@ I Introduction
Quantum sensors promise to estimate parameters with unprecedented precision, and are based on a mathematical primitive known as quantum state estimation. In quantum state estimation, the task is to estimate physical parameters embedded within quantum states with minimal error. The Cramér-Rao approach [1, 2, 3, 4, 5, 6], a prevalent technique in quantum state estimation which provides lower bounds on the minimum mean square error (MSE) of the estimate. Such lower bounds, known as Cramér-Rao bounds, use Fisher information obtained from quantum measurements. Since the Fisher information captures only the local structure of a statistical model, Cramér-Rao bounds are best suited for local estimation problems, where one assumes that the unknown parameter is within a small neighborhood of a known value. In multiparameter quantum state estimation, the Cramér-Rao approach is more complicated than in the single parameter case; unlike the single-parameter case, where the Cramér-Rao bound is simply the inverse of the quantum Fisher information [1], the multiparameter situation requires additional nontrivial techniques [6].
Despite the prevalence of Cramér-Rao bounds in quantum state estimation theory, one should take note that they are designed for the local estimation setting, and not the global estimation setting where neither the location nor the size of the parameters’ neighborhood are known [7, 8, 9, 10, 11, 12, 13]. Even with quantum state estimation problems use independent and identically distributed (IID) parametrized quantum states, the MSE for global estimation settings and the MSE for local estimation settings can differ. For example, in Ref. [7] which investigates the canonical phase estimation problem for qubit systems, it was shown that the MSE of the global estimation problem is $\pi^{2}$ times larger than the MSE of the local estimation problem. Similar observations were made in [12, Section 4], [13, Section 5]. Ref. [14] investigates the phase estimation problem using two bosonic modes, and similarly finds that the NOON state fails to saturate the Cramér-Rao in the global estimation setting while several studies [15, 16, 17, 18] implemented the NOON state for this aim according to the proposal [19, 20, 21].
@@
An even larger disparity between the MSE of global estimation and the MSE of local estimation problems has been observed [22, 23, 24, 25, 26], even in some classical parameter estimation problems that pertain to estimating entropy and Rényi entropy in classical systems in the IID setting. In the IID setting of estimating quantum entropy and quantum Renyi entropies of quantum states, [25, 26] similarly found that quantum Cramér-Rao bounds are not accurate for a global estimation strategy  [^1] This is because the Cramér-Rao bound is equal to the varentropy [27, 28], whose maximum value is upper bounded by $(\log d)^{2}$ [29, Lemma 8]. If the Cramér-Rao approach were to accurately describe the optimal performance of global estimation strategies with a constant error, then the number of copies of states needed (sample complexity) is $O(\log d)^{2}$ for a $d$-dimensional system. However, estimating both the classical and quantum entropy to a constant error requires much larger complexity, even in the classical setting [22, 23, 24, 25, 26]..
Given the possible disparity between the MSE between global and local estimation problems in the IID setting, one might not expect Cramér-Rao bounds to be accurate in a more complicated non-IID setting. However, it was surprisingly found that in the context of classical estimation theory that the non-IID situation globally estimating parametrised classical Markovian processes and classical hidden Markovian processes with a finite state system have MSE accurately described by Cramér-Rao bounds [30, 31].
Hence, it is non-trivial to determine whether a Cramér-Rao bound is accurate for global estimation problems, even in the IID setting. While the attainability of the Cramér-Rao bound has been studied in the one-copy setting [32], the attainability of Cramér-Rao bound for other non-IID settings has been less studied.
@@
Second, we proceed to families generated by unitary evolutions of $\mathop{\rm SU}(2)$ over probe states that begin in the number basis. Unlike many previous studies [7, 8, 9, 10, 40, 41, 42, 43, 44, 45, 46, 47, 11, 48, 49, 50] that focus on estimating mutiple phases in a similar setting, we specifically address the less studied scenario of the estimation of the state family generated by unitary evolutions of $\mathop{\rm SU}(2)$. The probe states can be in a (1) binomial distribution, (2) geometric distribution, and a (3) delta distribution in the number basis. For this, we consider the problem of estimating parameters embedded in a unitary model, where the unitary channel acts on the probe state. We analyze the global estimation of this problem by drawing an analogue between the bosonic system and the $\mathop{\rm SU}(2)$ system with a spin-$j$ system, and we employ the covariant approach initiated by Holevo [51], [2, Chapter 3]. This covariant approach allows us to solve a minimax problem with group symmetry [51], [2, Chapter 4], [52, Chapter 4]. Since we may describe global estimation using an appropriate minimax problem, the covariant approach works for the global estimation. By calculating both local and global estimation bounds, we are able to determine if the Cramér-Rao approach is accurate for the global estimation of our unitary model. For the half-Dicke state, we show that global estimation does not have the quantum advantage that local estimation promises.
The remaining part of this paper is organized as follows. First in Section II, we explain how bosonic states may arise in practice, and one may prepare the families of bosonic states that we consider in our paper. Second in Section III, we review the general formulation of quantum state estimation based on the Cramér-Rao approach. Third in Section IV, we discuss the attainability of the Cramér-Rao bound in the global estimation setting with respect to several quantum state estimation problems. Fourth in Section V, we discuss the attainability and the unattainability of the Cramér-Rao bound in the global estimation setting with respect to several quantum state estimation problems under a unitary model. Finally in Section VI, we have a final discussion of the results that we obtain.
II Bosonic states and their preparation
II.1 Boson Fock space and geometric distribution
@@
II.2 Symmetric space and binomial distribution
@@ II.3 Delta distribution
III General formulation of Cramér-Rao approach
@@
$C_{\bm{\theta}}[G]:=\min_{\hat{\Pi}\mathrm{\,:l.u.at\,}{\bm{\theta}}}\mathrm{Tr}{\big{[}GV_{\bm{\theta}}[\hat{\Pi}]\big{]}},$ (14)
@@
$C_{\bm{\theta}}[G]\geq C^{\rm R}_{\bm{\theta}}[G]:=\mathrm{Tr}\mathrm{Re}\,\sqrt{G}\tilde{F}^{-1}\sqrt{G}+\mathrm{Tr}|\mathrm{Im}\,\sqrt{G}\tilde{F}^{-1}\sqrt{G}|.$ (21)
@@
IV Attainability of the Cramér-Rao bound in the global estimation setting
@@
V Unattainability of the Cramér-Rao bound in the global estimation estimation setting
V.1 Local estimation of a unitary channel
@@
Then, given a parameter $\theta:=(\theta_{1},\theta_{2})$, we consider the state family $\rho_{\theta}:=U_{\theta}\rho U_{\theta}^{\dagger}$, where $U_{\theta}:=\exp(i(\theta_{1}J_{1}+\theta_{2}J_{2}))$. The two-parameter space $\Theta$ is given as $\{\theta|0\leq|\theta|\leq\pi\}$. This state family $\{\rho_{\theta}\}_{\theta}$ has two parameters, and is obtained from applying unitary operator $U_{\theta}$ on an initial probe state $\rho$. In this model, the SLD Fisher information is diagonal, in the sense that $F_{1,2}=F_{2,1}=0$. This implies that $C^{\rm S}_{\theta}[I]=F_{1,1}^{-1}+F_{2,2}^{-1}$. Furthermore,
@@
    Geometric distribution:- Consider $\rho$ as $\rho^{(n)}_{{\rm G},r}$. From Appendix A, this model satisfies the D-invariant condition. Hence, the RLD bound gives a tighter lower bound than the SLD bound. As calculated in Appendix C, when $r$ is fixed and $j=\frac{n}{2}$ increases, the RLD bound is approximated as
@@
V.2 A group covariant approach for global estimation
We consider the state family $\{\rho_{\theta}\}_{\theta}$ as given in Section V. Since the state family $\{\rho_{\theta}\}_{\theta}$ has a group covariant structure, we can employ a group covariant approach [51], [2, Chapter 4], [52, Chapter 4], where we employ a group covariant error function.
Now we consider the spin $j$ system ${\cal H}_{j}$ spanned by $\{|j;m\rangle\}_{m=-j}^{j}$. For an unknown value of $\theta$, and our estimate $\hat{\theta}$, we can denote the fidelity between the states $U_{\theta}|\frac{1}{2};\frac{1}{2}\rangle$ and $U_{\hat{\theta}}|\frac{1}{2};\frac{1}{2}\rangle$ as
@@ V.2 A group covariant approach for global estimation
When $\theta=0$, this fidelity simplifies to $R(0,\hat{\theta})=\cos^{2}(\hat{\theta}/2).$ We define the error function of our estimate to be $\eta(\hat{\theta})\coloneqq 4(1-R(0,\hat{\theta}))$, which to leading order in $\hat{\theta}$ is equivalent to $|\hat{\theta}|^{2}$. In this definition, the error function $\eta(\hat{\theta})$ is to leading order in $\hat{\theta}$ equivalent to the MSE of $\hat{\theta}$ in the local estimation scenario when $\theta=0$.
@@
While the function $R(M)$ is more difficult to calculate than $R_{\nu}(M)$, the situation simplifies greatly when the optimal POVM $M$ is covariant. In this situaton, the Bayesian error function can be equal to the worst-case error function in the sense that $1-R_{\nu}(M)=1-R(M)$ [51], [2, Chapter 4], [52, Chapter 4]. This situation is possible when $R(g\theta,g\hat{\theta})=R(\theta,\hat{\theta})$ and when $\nu$ is invariant under any $g$. In our scenario the POVM that maximizes $R_{\nu}(M)$ ($R(M)$) is realized by a covariant POVM [51], [2, Chapter 4], [52, Chapter 4]. In such a situation, we can calculate $R_{\nu}(M)$ instead of $R_{\nu}(M)$.
@@
$\frac{1}{2j+2}\mathrm{Tr}P_{j+\frac{1}{2}}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho\geq\frac{1}{2j}\mathrm{Tr}P_{j-\frac{1}{2}}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho$ (47)
@@
For reader’s convenience, we give a proof for Theorem 1 in Appendix D as a special case of [52, Theorem 4.6]. The proof of Theorem 1 uses the following ideas. First, we represent the error function using a spin-1/2 representation while considering the estimation of a quantum state in the spin-$j$ representation. Second, we apply the definition of $R(M)$ and use the fact that the product of traces is equal to the trace of the tensor products of the arguments, and use Schur’s lemma appropriately.
V.3 Global estimation of a unitary channel
Here we apply the theory reviewed in Section V.2 to calculate the minimum MSE for the global estimation of our unitary model.
V.3.1 Probe state as a binomial distribution in the number basis

### v5

@@ 1 Introduction
Quantum sensors promise to estimate parameters with unprecedented precision, and are based on a mathematical primitive known as quantum state estimation. In quantum state estimation, the task is to estimate physical parameters embedded within quantum states with minimal error. The Cramér-Rao approach [1, 2, 3, 4, 5, 6], a prevalent technique in quantum state estimation which provides lower bounds on the minimum mean square error (MSE) of the estimate. Such lower bounds, known as Cramér-Rao bounds (CRBs), use Fisher information obtained from quantum measurements. Since the Fisher information captures only the local structure of a statistical model, Cramér-Rao bounds are best suited for local estimation problems, where one assumes that the unknown parameter is within a small neighborhood of a known value. In multiparameter quantum state estimation, the Cramér-Rao approach is more complicated than in the single parameter case; unlike the single-parameter case, where the CRB is simply the inverse of the quantum Fisher information [1], the multiparameter situation requires additional nontrivial techniques [6].
Despite the prevalence of CRBs in quantum state estimation theory, one should note that they are fundamentally designed for local estimation settings, where the parameter’s neighborhood is known. They are not directly formulated for global estimation settings, where the true parameter’s location is unknown and the entire parameter space must be considered [7, 8, 9, 10, 11, 12, 13]. While the MSE for global estimation is naturally lower bounded by the MSE for local estimation, and thus by a CRB, a CRB truly represents the precision of an estimator only if the MSE for global estimation actually achieves that CRB.
To justify this attainability, many researchers traditionally consider the asymptotic setting with independent and identically distributed (IID) quantum states. Indeed, studies like Ref. [4] have shown that a specific type of CRB (the Holevo type) can be asymptotically attained for global estimation in the IID setting, even for multiple parameters, when the model is fixed. Furthermore, if a state family forms an exponential family, such as thermal states [14, 15], the CRB can be attained without requiring an anymptotic limit [16, Theorem 6.7][17, 18].
The success of CRBs in these specific state estimation scenarios has led to a widespread expectation that the CRB can also be attained in channel estimation when multiple uses of the same unknown channel are available. This situation is often implicitly treated as an “IID setting” for channels within the community. For example, in the canonical phase estimation problem for qubit systems (an estimation of unitary), Refs. [19, 20, 21] focused on this problem, demonstrating that the maximum Fisher information achieves Heisenberg scaling and showing that this maximum is attained when the NOON state is used as input. (The NOON state has been experimentally implemented by several studies [22, 23, 24, 25].) Building on these findings, they concluded that the phase can be estimated with Heisenberg scaling.
However, this assumption that the CRB is always attainable under ”IID” conditions, particularly in global estimation, is often inaccurate. Our paper aims to highlight this critical distinction. For instance, in the canonical phase estimation problem for qubit systems, Ref. [7] showed that the MSE of the global estimation problem is $\pi^{2}$ times larger than the MSE of the local estimation problem, whose performance matches Fisher information. Similar discrepancies were observed in Ref. [12, Section 4] and Ref. [13, Section 5]. Furthermore, Ref. [26] found that the NOON state similarly fails to saturate the CRB in the global estimation setting for phase estimation using two bosonic modes. These facts demonstrate the significant difficulty in establishing a direct relationship between the MSE for global estimation and CRBs in channel estimation.
@@
It is crucial to note that even the IID setting for state estimation is not universally sufficient to guarantee CRB attainability for global estimation. When the complexity of the state family (i.e., the model) increases with the number of available copies, CRBs generally cannot be saturated. For example, in classical systems, CRBs cannot be saturated for estimating entropy and Rényi entropy in the IID setting when the system size increases [27, 28, 29]. Similarly, for quantum entropy and quantum Rényi entropies, quantum CRBs are found to be inaccurate for global estimation strategies when the system size increases [30, 31, 32, 33]  [^1] This is because the CRB in these cases equals the varentropy [32, 33], which is upper bounded by $(\log d)^{2}$ [34, Lemma 8]. If the CRB accurately described optimal global estimation performance with a constant error, the sample complexity would be $O(\log d)^{2}$ for a $d$-dimensional system. However, estimating both classical and quantum entropy to a constant error requires much larger complexity, even in the classical setting [27, 28, 29, 30, 31].
Despite these challenges, it has been shown that the CRB can still be achieved in global estimation settings for channel estimation when the maximum Fisher information increases linearly with respect to the number of uses of the channel [35]  [^2] The idea in Ref. [35] is to consider $m$ uses of a channel as a single effective channel, prepare an optimal input state for it, and then repeat this process $k$ times, viewing it as a state estimation problem. Under these conditions, the CRB can be attained if the maximum Fisher information scales linearly with the number of uses of the channel. Recently, Ref. [36] employed this idea to achieve what they term Heisenberg scaling. However, in this case, the inverse of the MSE behaves as $O(km^{2})$ while the total number of channel uses is $km$, meaning that true Heisenberg scaling, which demands an $O((km)^{2})$ scaling, is not achieved. Given the possible disparity between the MSE for global and local estimation problems even in the IID setting, one might not expect CRBs to be accurate in more complicated non-IID settings. Surprisingly, however, in classical estimation theory, it was found that for non-IID situations, such as globally estimating parametrized classical Markovian processes and classical hidden Markovian processes with a finite state system, the MSE is accurately described by CRBs [37, 38]. Nevertheless, the inherent difficulties in guaranteeing CRB attainability have led some researchers to employ complementary approaches like Ziv-Zakai Error Bounds, which offer alternative perspectives on estimation limits [39, 40]. Furthermore, as was known even in the physics community [41], the CRB cannot be attained with finite samples, even in the classical IID setting with a fixed model.
Initial steps tackling the attainability of the CRB have been taken. For example, Ref. [42, 43] described a numerical approach based on semidefinite programs that calculates global estimation bounds from local estimation bounds on a fictitious state. Moreover, the attainability of the CRB in the one-copy setting has been studied [16, Theorem 6.7][17, 18, 44]. However, the attainability of the CRB for other non-IID settings has been less studied. Indeed, determining whether a CRB accurately describes global estimation problems remains a non-trivial challenge, even in the IID setting.
@@
Second, we proceed to families generated by unitary evolutions of $\mathop{\rm SU}(2)$ over probe states that begin in the number basis. Unlike many previous studies [7, 8, 9, 10, 52, 53, 54, 55, 56, 57, 58, 59, 11, 60, 61, 62] that focus on estimating mutiple phases in a similar setting, we specifically address the less studied scenario of the estimation of the state family generated by unitary evolutions of $\mathop{\rm SU}(2)$. The probe states can be in a (1) binomial distribution, (2) geometric distribution, and a (3) delta distribution in the number basis. For this, we consider the problem of estimating parameters embedded in a unitary model, where the unitary channel acts on the probe state. We analyze the global estimation of this problem by drawing an analogue between the bosonic system and the $\mathop{\rm SU}(2)$ system with a spin-$j$ system, and we employ the covariant approach initiated by Holevo [63], [2, Chapter 3]. Fortunately, this covariant approach guarantees that a covariant estimator achieves the optimal estimator with group symmetry [63], [2, Chapter 4], [64, Chapter 4], which allows us to restrict our estimator within covariant estimators. Since we may describe global estimation using an appropriate minimax problem, the covariant approach works for the global estimation. By calculating both local and global estimation bounds, we are able to determine if the Cramér-Rao approach is accurate for the global estimation of our unitary model. For the half-Dicke state, we show that global estimation does not have the quantum advantage that local estimation promises.
The remaining part of this paper is organized as follows. First in Section 2, we explain how bosonic states may arise in practice, and one may prepare the families of bosonic states that we consider in our paper. Second in Section 3, we review the general formulation of quantum state estimation based on the Cramér-Rao approach. Third in Section 4, we discuss the attainability of the Cramér-Rao bound in the global estimation setting with respect to several quantum state estimation problems. Fourth in Section 5, we discuss the attainability and the unattainability of the Cramér-Rao bound in the global estimation setting with respect to several quantum state estimation problems under a unitary model. Finally in Section 6, we have a final discussion of the results that we obtain.
We reiterate that our paper allows one to calculate the optimal global estimation bound for a unitary channel that acts on a symmetric probe state, and we show that such a global bound need not be equal to the local estimation bound from the Cramér-Rao approach.
2 Bosonic states and their preparation
2.1 Boson Fock space and geometric distribution
@@
2.2 Symmetric space and binomial distribution
@@ 2.3 Delta distribution
3 General formulation of Cramér-Rao approach
@@
$C_{\bm{\theta}}[G]:=\min_{\hat{\Pi}\mathrm{\,:l.u.at\,}{\bm{\theta}}}\mathrm{Tr}{\big{[}GV_{\bm{\theta}}[\hat{\Pi}]\big{]}},$
@@
$C_{\bm{\theta}}[G]\geq C^{\rm R}_{\bm{\theta}}[G]:=$ $\mathrm{Tr}\mathrm{Re}\,\sqrt{G}\tilde{F}^{-1}\sqrt{G}$
$+\mathrm{Tr}|\mathrm{Im}\,\sqrt{G}\tilde{F}^{-1}\sqrt{G}|.$ (20)
@@
4 Attainability of the Cramér-Rao bound in the global estimation setting
@@
5 Unattainability of the Cramér-Rao bound in the global estimation estimation setting
5.1 Local estimation of a unitary channel
@@
Then, given a parameter $\theta:=(\theta_{1},\theta_{2})$, we consider the state family $\rho_{\theta}:=U_{\theta}\rho U_{\theta}^{\dagger}$, where $U_{\theta}:=\exp(i(\theta_{1}J_{1}+\theta_{2}J_{2}))$. The two-parameter space $\Theta$ is given as $\{\theta||\theta|\leq\pi\}$. This state family $\{\rho_{\theta}\}_{\theta}$ has two parameters, and is obtained from applying unitary operator $U_{\theta}$ on an initial probe state $\rho$. In this model, the SLD Fisher information is diagonal, in the sense that $F_{1,2}=F_{2,1}=0$. This implies that $C^{\rm S}_{\theta}[I]=F_{1,1}^{-1}+F_{2,2}^{-1}$. Furthermore,
@@
    Geometric distribution:- Consider $\rho$ as $\rho^{(n)}_{{\rm G},r}$. From Appendix B, this model satisfies the D-invariant condition. Hence, the RLD bound gives a tighter lower bound than the SLD bound. As calculated in Appendix D, when $r$ is fixed and $j=\frac{n}{2}$ increases, the RLD bound is approximated as
@@
5.2 A group covariant approach for global estimation
We consider the state family $\{\rho_{\theta}\}_{\theta}$ as given in Section 5. Since the state family $\{\rho_{\theta}\}_{\theta}$ has a group covariant structure, we can employ a group covariant approach [63], [2, Chapter 4], [64, Chapter 4], where we employ a group covariant error function. This is because a covariant POVM realizes the optimal performance under the symmetric setting [63], [2, Chapter 4], [64, Chapter 4].
Now we consider the spin $j$ system ${\cal H}_{j}$ spanned by $\{|j;m\rangle\}_{m=-j}^{j}$. For an unknown value of $\theta$, and our estimate $\hat{\theta}$, we employ the fidelity between the states $U_{\theta}|\frac{1}{2};\frac{1}{2}\rangle$ and $U_{\hat{\theta}}|\frac{1}{2};\frac{1}{2}\rangle$ as
@@ 5.2 A group covariant approach for global estimation
Using the parameter $\phi$ as $e^{-i\phi}|\theta|=\theta_{1}+i\theta_{2}$, as shown in Appendix A, the fidelity is calculated as
$R(\theta,\hat{\theta})=$ $\Big{|}\cos\frac{|\theta|}{2}\cos\frac{|\hat{\theta}|}{2}+e^{i(-\phi+\hat{\phi})}\sin\frac{|\theta|}{2}\sin\frac{|\hat{\theta}|}{2}\Big{|}^{2}.$ (41)
In particular, when $\theta=0$, this fidelity simplifies to $R(0,\hat{\theta})=\cos^{2}(|\hat{\theta}|/2)$. Then, given a parameter $\theta:=(\theta_{1},\theta_{2})$, we consider the state family $\rho_{\theta}:=U_{\theta}\rho U_{\theta}^{\dagger}$.
We define the error function of our estimate to be $\eta(\hat{\theta})\coloneqq 4(1-R(0,\hat{\theta}))$, which to leading order in $\hat{\theta}$ is equivalent to $|\hat{\theta}|^{2}$. In this definition, the error function $\eta(\hat{\theta})$ is to leading order in $\hat{\theta}$ equivalent to the MSE of $\hat{\theta}$ in the local estimation scenario when $\theta=0$. Since the fidelity is unitarily invariant, the error function $\eta(\hat{\theta})$ satisfies the group covariance condition required for the group covariant approach.
@@
While the function $R(M)$ is more difficult to calculate than $R_{\nu}(M)$, the situation simplifies greatly when the optimal POVM $M$ is covariant. In this situation, the Bayesian error function can be equal to the worst-case error function in the sense that $1-R_{\nu}(M)=1-R(M)$ [63], [2, Chapter 4], [64, Chapter 4]. This situation is possible when $R(g\theta,g\hat{\theta})=R(\theta,\hat{\theta})$ and when $\nu$ is invariant under any $g$. In our scenario the POVM that maximizes $R_{\nu}(M)$ ($R(M)$) is realized by a covariant POVM [63], [2, Chapter 4], [64, Chapter 4]. In such a situation, we can calculate $R(M)$ as well as $R_{\nu}(M)$.
@@
$\frac{1}{2j+2}\mathrm{Tr}P_{j+\frac{1}{2}}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho$
$\geq$ $\frac{1}{2j}\mathrm{Tr}P_{j-\frac{1}{2}}|\frac{1}{2};\frac{1}{2}\rangle\langle\frac{1}{2};\frac{1}{2}|\otimes\rho$ (47)
@@
For reader’s convenience, we give a proof for Theorem 1 in Appendix E as a special case of [64, Theorem 4.6]. The proof of Theorem 1 uses the following ideas. First, we represent the error function using a spin-1/2 representation while considering the estimation of a quantum state in the spin-$j$ representation. Second, we apply the definition of $R(M)$ and use the fact that the product of traces is equal to the trace of the tensor products of the arguments, and use Schur’s lemma appropriately.
5.3 Global estimation of a unitary channel
Here we apply the theory reviewed in Section 5.2 to calculate the minimum MSE for the global estimation of our unitary model.
5.3.1 Probe state as a binomial distribution in the number basis